Diane has $1.70 in dimes and nickels. She has a total of 22 coins. How many of each kind does she have?
Nickels
Dimes

Answer by Mathtut(3670)

let d and n be the number of dimes and nickels respectively
:
d+n=22..........eq 1
.1d+.05n=1.70...eq 2
:
rewrite eq 1 to d=22-n and plug this value into eq 2
:
.1(22-n)+.05n=1.7
:
2.2-.1n-+.05n=1.7
:
-.05n=-.5
:
n=10 nickels
:
d=22-n=22-10=12 dimes